# DepthSplat: Connecting Gaussian Splatting and Depth

### v1

@@ 4.4 Benchmark Comparisons
Figure 7: Visual Comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat (Chen et al., 2024) on regions that are hard to match (e.g., repeated patterns).
@@ 5 Conclusion
In this paper, we introduce DepthSplat, a new approach to connecting Gaussian splatting and depth to achieve state-of-the-art results on ScanNet, RealEstate10K and DL3DV datasets for both depth and view synthesis tasks. We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabelled multi-view image datasets for training more multi-view consistent and robust depth models. Our current model requires camera pose information as input along with the multi-view images, removing this requirement would be exciting future work.
Acknowledgement
@@
Appendix A Appendix
A.1 Depth Pre-Training for Gaussian Splatting
We further study the effect of depth pre-training for Gaussian splatting experiments. Unlike the pre-trained models Depth Anything (Yang et al., 2024b) and UniMatch (Xu et al., 2023), which are trained for either monocular and multi-view features separately, we perform joint training of our full two-branch depth model on the depth datasets. We then compare the results of different initializations to the depth network for the full Gaussian splatting model. We can see from Table 7 that improved depth initialization leads to better view synthesis results.
Table 7: Depth to Gaussian Splatting Transfer. We compare different pre-trained network weights for initializing the depth network when training our full DepthSplat model for view synthesis. Compared to using Depth Anything and UniMatch pre-trained monocular and multi-view network weights for initializing the monocular ViT and multi-view Transformer features, our jointly trained 2-branch depth model (full model) performs best on all metrics.
@@ A.1 Depth Pre-Training for Gaussian Splatting
Depth Anything (only mono) 26.59 0.874 0.1256
UniMatch + Depth Anything (mv & mono) 26.76 0.877 0.1234
DepthSplat (full model) 26.81 0.878 0.1225
A.2 More Visual Comparisons on DL3DV
We show more visual comparison results on DL3DV in Fig. 8 with different number of input views, where our DepthSplat consistently outperforms MVSplat in challenging regions.
Figure 8: Different number of input views on DL3DV. Our DepthSplat performs consistently better than MVSplat (Chen et al., 2024).

### v2

@@ 4.4 Benchmark Comparisons
Figure 7: Visual Comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat [8] on regions that hard to match (e.g., repeated patterns).
Comparisons on DL3DV. To evaluate the performance on complex real-world scenes, we conduct comparisons with the representative model MVSplat [8] on the recently introduced DL3DV dataset [30]. We also compare the results of different numbers of input views (2, 4 and 6) on this dataset. We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Tab. 6. Our DepthSplat consistently outperforms MVSplat in all metrics, and our improvements are more significant with fewer input views, which indicates that our model is more robust to sparse views. Visual comparisons with MVSplat on the DL3DV dataset are shown in Fig. 7, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results. We show more visual comparison results in Fig. C.3 of the appendix. It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec. 3.1), which is unlike the expensive global pair-wise matching used in MVSplat.
Table 7: Cross-dataset generalization. When generalizing to DL3DV with RealEstate10K pre-trained models, our DepthSplat is consistently better than MVSplat for both small and large baselines.
Method Small Baseline Large Baseline
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MVSplat [8] 19.19 0.608 0.304 16.26 0.459 0.438
DepthSplat 20.29 0.672 0.259 17.50 0.536 0.376
Cross-dataset generalization. We further evaluate the cross-dataset generalization capability of our model by directly testing on DL3DV with RealEstate10K pre-trained models. The results are shown in Tab. 7, where our DepthSplat consistently outperforms MVSplat for both small and large baseline inputs (visual comparisons are shown in Fig. C.2 of the appendix). We also show some visual results of our depth model on unseen samples in Fig. 6, where our model generalizes robustly to different scene types.
High-resolution results. In addition to the comparison results on the $256\times 256$ resolution RealEstate10K and $256\times 448$ DL3DV dataets, we show more high-resolution ($512\times 960$) qualitative results in Fig. C.4 and Fig. C.5 of the appendix. We also invite the readers to our project page haofeixu.github.io/depthsplat for high-resolution ($512\times 960$) video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views.
@@ 5 Conclusion
In this paper, we introduce DepthSplat, a new approach to connecting Gaussian splatting and depth to achieve state-of-the-art results on ScanNet, RealEstate10K and DL3DV datasets for both depth and view synthesis tasks. We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabeled multi-view image datasets for training more robust and generalizable multi-view depth models. Our current model requires camera pose information as input along with the multi-view images, removing this requirement would be exciting future work.
6 Acknowledgement
@@
Appendix
A Model Efficiency
We report the number of parameters, inference time and memory consumption in Tab. A.1 for different model sizes. In particular, the “small” model refers to “ViT-S monocular, 1-scale multi-view” in Tab. 1 of the main paper, and the “base” model refers to “ViT-B monocular, 2-scale multi-view”, and the “large” model refers to “ViT-L monocular, 2-scale multi-view”. We report the inference time (measured on an A100 GPU) and memory for 2 input views at $256\times 256$ and $512\times 512$ resolutions. Our DepthSplat can support different speed-accuracy trade-offs.
Table A.1: Model efficiency.
Model Size Resolution Params (M) Time (s) Memory (GB)
Small $256\times 256$ 37 0.065 1.4
Base 117 0.081 2.1
Large 360 0.105 3.5
Small $512\times 512$ 37 0.091 3.1
Base 117 0.155 5.4
Large 360 0.313 8.4
B Depth Pre-Training for Gaussian Splatting
In Sec. 4.1 of the main paper, we have shown that better depth architectures lead to improved view synthesis with Gaussian splatting. We further study the effect of different weight initializations for the depth model. Specifically, we compare three variants: 1) only initializing the monocular feature with Depth Anything V2 [66]; 2) initializing the monocular feature with Depth Anything V2 [66] and the multi-view feature with UniMatch [61]; 3) initializing the full depth model by pre-training the depth model on TartanAir [55]. We can observe from Tab. B.1 that better depth initialization also leads to improved view synthesis results.
Table B.1: Depth pre-training for Gaussian splatting. We compare different weight initializations for the depth model when training our full DepthSplat model for view synthesis. Compared to 1) only initializing the monocular feature (mono features) and 2) initializing the monocular and multi-view features (mono & mv features), our pre-trained full depth model (full depth model) achieves the best view synthesis results.
@@ B Depth Pre-Training for Gaussian Splatting
mono features 26.59 0.874 0.1256
mono & mv features 26.76 0.877 0.1234
full depth model 26.81 0.878 0.1225
C More Visual Results
C.1 Unsupervised Depth Pre-Training with Gaussian Splatting
In Fig. C.1, we show the comparisons of depth estimation results with and without Gaussian splatting pre-training. We observe that pre-training leads to better results in texture-less regions. We hypothesize that pre-training provides regularizations (as also observed in  [19]) to the challenging scenarios and accordingly leads to improved performance.
Figure C.1: Effect of unsupervised depth pre-training. We observe that the unsupervised pre-training leads to improved performance for texture-less regions.
C.2 Cross-Dataset Generalization
In Fig. C.2, we show the cross-dataset generalization results on the DL3DV [30] dataset, which are obtained with the RealEstate10K [75] trained models. Our DepthSplat generalizes more robustly than MVSplat [8] on unseen scenes.
Figure C.2: Generalization from RealEstate10K to DL3DV.
C.3 Visual Comparisons on DL3DV
In Fig. C.3, we compare the visual synthesis results from 4 input views with MVSplat [8]. Our DepthSplat better preserves the scene structures.
Figure C.3: View synthesis from 4 input views on DL3DV.
C.4 High-Resolution Results
In Fig. C.4 and Fig. C.5, we show the view synthesis and multi-view depth estimation results at $512\times 960$ resolutions. Note that MVSplat [8] runs out-of-memory on such high resolutions with 6 or 12 input views, while our DepthSplat significantly improves the efficiency with two technical components. First, our lowest feature resolution is $1/8$ of the image resolution, while MVSplat uses $1/4$. Second, we use local cross-view attentions (detailed in Sec. 3.1 of the main paper) unlike the pair-wise global attentions in MVSplat.
Figure C.4: View Synthesis at $512\times 960$ resolutions from 6 input views.
Figure C.5: Depth predictions on DL3DV with 12 input views. The image resolutions are $512\times 960$.
D More Implementation Details
We provide more implementation details on high-resolution experiments. For high-resolution experiments, we choose our small model which contains a ViT-S monocular branch and a single-scale multi-view branch. We first train our model on RealEstate10K at $448\times 768$ resolutions for 200K iterations with a batch size of 8, and then we fine-tune the model on DL3DV at $448\times 768$ resolutions for 200K iterations with a batch size of 8. We use random number of input views (2 to 6) for training. Despite that our model is trained at $448\times 768$ resolutions with at most 6 input views, we observe our model can generalize reasonably well on higher resolutions (e.g., $512\times 960$) and more input views (e.g., 12), as we show in our project page: haofeixu.github.io/depthsplat.